# Tool-to-Agent Retrieval: Bridging Tools and Agents for Scalable LLM Multi-Agent Systems

### v1

@@ 3 Experiments and Evaluation
We evaluate retrieval performance across several embedding models. Specifically, we use Vertex AI embeddings from Google [8], Titan embeddings from Amazon [1], and OpenAI embedding models [26]. The datasets are embedded using each model, and semantic similarity search is performed to retrieve relevant entities. We first retrieve the top $N\gg K$ entities from the tool–agent catalog and then select the top-$K$ unique agents using Algorithm 1. Retrieval accuracy is computed by comparing the retrieved agents against the ground-truth agents associated with each query in the evaluation set.
Approach Recall mAP nDCG
@1 @3 @5 @1 @3 @5 @1 @3 @5
BM25 0.205 0.20 0.20 0.12 0.12 0.12 0.14 0.14 0.14
Q.Retrieval 0.31 0.47 0.56 0.31 0.31 0.24 0.310 0.35 0.32
MCPZero 0.44 0.66 0.70 0.45 0.39 0.31 0.45 0.46 0.41
ScaleMCP 0.49 0.68 0.74 0.49 0.40 0.29 0.49 0.48 0.40
RRF 0.58 0.74 0.79 0.58 0.44 0.33 0.59 0.52 0.44
Tool-to-Agent Retrieval 0.61 0.77 0.83 0.61 0.49 0.34 0.61 0.56 0.46
@@ Results.
Vertex AI text-embedding-005 0.87 0.74 0.48 0.42 0.36 0.32
Gemini-embedding-001 0.86 0.74 0.49 0.44 0.37 0.34
Amazon titan-embed-text-v2 0.85 0.66 0.47 0.37 0.35 0.28
Amazon titan-embed-text-v1 0.85 0.65 0.48 0.39 0.36 0.30
OpenAI text-embedding-ada-002 0.83 0.70 0.50 0.40 0.39 0.30
OpenAI text-embedding-3-small 0.87 0.72 0.49 0.41 0.36 0.31
OpenAI text-embedding-3-large 0.87 0.74 0.50 0.42 0.38 0.32
All-MiniLM-L6-v2 0.80 0.67 0.45 0.39 0.33 0.29

### v2

@@ 2 Approach: Tool-to-Agent Retrieval
Approach Recall mAP nDCG
@1 @3 @5 @1 @3 @5 @1 @3 @5
BM25 0.20 0.20 0.20 0.12 0.12 0.12 0.14 0.14 0.14
Q.Retrieval 0.31 0.47 0.56 0.31 0.31 0.24 0.31 0.35 0.32
MCPZero 0.44 0.66 0.70 0.45 0.39 0.31 0.45 0.46 0.41
ScaleMCP 0.49 0.68 0.74 0.49 0.40 0.29 0.49 0.48 0.40
Tool-to-Agent Retrieval 0.61 0.77 0.83 0.61 0.49 0.34 0.61 0.56 0.46
@@ 3 Experiments and Evaluation
We evaluate retrieval performance across several embedding models. Specifically, we use 8 embedding models, both closed and open source [8, 1, 26]. The datasets are embedded using each model, and semantic similarity search is performed to retrieve relevant entities. We first retrieve the top $N\gg K$ entities from the tool–agent catalog and then select the top-$K$ unique agents using Algorithm 1. Retrieval accuracy is computed by comparing the retrieved agents against the ground-truth agents associated with each query in the evaluation set.
@@ Results.
Vertex AI text embedding 005 0.87 0.74 0.48 0.42 0.36 0.32
Gemini Embedding 001 0.86 0.74 0.49 0.44 0.37 0.34
Amazon titan embed text v2 0.85 0.66 0.47 0.37 0.35 0.28
Amazon titan embed text v1 0.85 0.65 0.48 0.39 0.36 0.30
OpenAI text embedding ada 002 0.83 0.70 0.50 0.40 0.39 0.30
OpenAI text embedding 3 small 0.87 0.72 0.49 0.41 0.36 0.31
OpenAI text embedding 3 large 0.87 0.74 0.50 0.42 0.38 0.32
All MiniLM L6 v2 0.80 0.67 0.45 0.39 0.33 0.29